# Normalizing flows as an enhanced sampling method for atomistic supercooled liquids

### v1

@@ I Introduction
One of the most important methodological revolution in science in the last century is scientific computing [1]. Numerical simulations represent a way, complementary to experiments, to study physical systems, thus providing a unique lens on the microscopic mechanisms underpinning macroscopic physical phenomena [1, 2, 3, 4].
@@
A major application of numerical simulations, from the very beginning, has been sampling physical configurations at thermal equilibrium [5]. This was initially done by using either Monte Carlo Markov chains [5, 6] or molecular dynamics [7, 8]. Both methods can be viewed as ways to implement some physical dynamics to ergodically explore the configuration space, just as the physical system does. The basic challenge is to run those dynamics long enough to be able to generate a large set of uncorrelated configurations to perform accurate ensemble averages of physical observables [2].
@@
When the system is characterized by large relaxation times, for instance near phase transitions or in disordered media, sampling can become so challenging that conventional methods may fail [2, 4]. In such cases, the only solution, so far, consists in devising alternative dynamics that ensure equilibrium sampling while being characterised by substantially smaller decorrelation times. Such strategies are described in many standard textbooks on computer simulations and statistical physics [4, 3, 9, 10]. In the context of off-lattice molecular simulations, we can mention non-local [11, 12, 13], lifting [14], or collective [15, 16, 17] Monte Carlo algorithms, parallel tempering [18, 19, 20, 21], population annealing [22, 23, 24], or irreversible Langevin dynamics [25].
@@
However, there exist physical systems in which equilibrium sampling remains a difficult open problem even with enhanced sampling methods. Among them, glassy systems [26] stand as one of the most difficult to simulate in condensed and soft-matter physics. Molecular, colloidal and spin glasses are in fact known to display an extremely slow physical dynamics, which creates a major challenge to standard simulation algorithms [27, 28]. Glassy systems can in fact serve as a severe test of any newly proposed method, and can be seen as a paradigm for complex systems.
@@
The recent discovery of generative models in artificial intelligence (AI) able to generate large structured data such as images, sound, 3D-video, and text has the potential to induce a second revolution in scientific computing [29, 30, 31, 32, 33]. These AI models are not only able to accurately produce complex data, but are also very fast. Speed is a central requirement in the realm of scientific computing. Several applications appeared already. In 2019, Noé et al. proposed the usage of normalizing flows (NF) [34], and independently Wu et al. variational autoregressive models [35], for Boltzmann sampling in statistical and condensed matter physics. These works have found numerous applications for sampling [36, 37, 38, 39, 40] and free energy calculations [41, 42]. However, despite these interesting premises, a clear view on when, where and how these methods work, and in particular their limitations and efficiency against known algorithms, is currently lacking. For standard phase transitions, promising results have been obtained in [43, 44]. However, for hard computational problems such as complex and glassy systems, it is unclear whether they can circumvent the problem of large relaxation times. Positive results have been reported in [45, 46] for spin-glasses. On the other hand, theoretical and numerical analysis of mean-field models for structural glasses [47, 48], related to other hard problems in computer science [49], has shown that several generative models do not, and sometimes can not, have good performances. Worse, they sometimes perform less efficiently than conventional algorithms, such as local Monte Carlo [47]. These results provide a rather pessimistic view of the potentiality of machine learning (ML) techniques in the field of glassy systems. A last difficulty is that the performance of generative models is generically expected to scale very badly with the size of the system, so that applications to study phase transitions and collective effects in many-body systems appear out of sight.
@@
Given the rapid progress made in ML studies [50, 51, 52, 53, 54], we feel that there is room for hope and progress. At the moment, there is a clear need of further studies to develop and test generative models in hard computational problems, and benchmark their performances against the ones of existing algorithms. The aim of this work is to perform such analysis for atomistic models of glass-forming liquids [27]. Contrary to [47, 48], we study an off-lattice, finite-dimensional glassy model which displays an extremely slow dynamics, associated to a super-Arrhenius evolution of relaxation and sampling times. It is challenging to perform numerical simulations in realistic experimental conditions for this model, as the physical relaxation time increases by more than 14 orders of magnitude towards the experimental glass transition temperature [55].
@@
We focus on a specific two-dimensional glass-forming model that shows conventional signatures of glassy dynamics [56, 57], and represents therefore a relevant and challenging test bench for enhanced sampling methods. At low temperatures, molecular dynamics becomes rapidly unable to perform an equilibrium sampling of the configuration space, even for modest system sizes. Given the challenges mentioned above, we intentionally study a relatively small system size to separate the capabilities of NF to tackle complex landscapes from its potentially problematic scaling with system size.
@@
We optimise a ML technique based on normalizing flows [33, 34], which we carefully benchmark against several advanced techniques introduced in the physics literature, such as parallel tempering [19], population annealing [23], and swap Monte Carlo [58]. By studying and comparing their abilities to produce an ensemble of equilibrated low-temperature configurations, we provide the first quantitative analysis of the performance of ML methods to sample realistic supercooled liquids at low temperatures. Surprisingly, our results demonstrate the great potential of such method which turns out to be much more efficient than conventional molecular dynamics and achieves performances comparable to parallel tempering and population annealing. Finally, we assess current limitations of these new methods and provide guidelines for further studies, in particular to improve the parametrization of the normalizing flow and to extend this technique to larger system sizes.
@@ II.1 A two-dimensional ternary Lennard-Jones mixture
We study a two-dimensional model introduced and developed in Ref. [56]. This is a variation of the binary Lennard-Jones mixture introduced long ago by Kob and Andersen [59] in which a third component is introduced to both improve the glass-forming ability (i.e. to prevent easy crystallization) and enable a more efficient use of the swap Monte Carlo algorithm, a strategy proposed in [58, 60]. We refer to [56, 57] for all details regarding the model parameters and simulation details.
We study systems with $N=43$ particles using periodic boundary conditions with box length $L=6.0$, in reduced units. The unit of length is $\sigma$, which corresponds to the diameter of the large particles. When using molecular dynamics, the unit of time is the Lennard-Jones timescale $\tau=\sqrt{m\sigma^{2}/\epsilon}$ where $m$ is the particle mass, and $\epsilon$ the interaction strength between large particles. For different algorithms, we express times in units of $\tau$, in order to carefully reflect the actual computational cost of each method.
The relatively small system size is actually comparable to previous studies using normalizing flows for sampling in complex systems [34, 36, 39]. Results for a larger system size, $N=172$, are presented in Appendix C. The main goal of this work is to benchmark the efficiency of normalizing flows in small glassy systems. Scaling to larger systems introduces additional challenges that have to be considered separately and are left for future work. While $N=43$ seems a small number of particles, we emphasize that this is large enough to produce glassy dynamics and local structure for this dense fluid that are nearly equivalent to those of much larger systems [61]. This implies in particular that equilibrium sampling even for this modest system size is already a difficult computational challenge.
@@ II.2 The specific heat as a sampling task
To solve this problem we analyse the statistics of energy fluctuations measured in an ensemble of configurations. For each configuration, we measure the total potential energy, $E_{\rm pot}=\sum_{i\neq j}V(r_{ij})$, where $V(r)$ denotes the pair interaction potential, and $r_{ij}$ the relative distance between particles $i$ and $j$. We define the average potential energy over this ensemble, $\langle E_{\rm pot}\rangle$, and the variance of its fluctuations, which is directly connected to the specific heat as
$c_{V}=\frac{C_{V}}{N}=\frac{\langle E_{\rm pot}^{2}\rangle-\langle E_{\rm pot}\rangle^{2}}{NT^{2}}.$ (1)
Empirically, we find that a correct estimate of the specific heat at a given temperature requires the production of about 100 independent equilibrium configurations. We conclude that the determination of the specific heat represents a well-defined task that is able to probe the capability of a given algorithm to (i) reach thermal equilibrium, (ii) sample a large number of independent configurations representative of the Boltzmann distribution. Another advantage is that $c_{V}$ does not require knowledge of a dynamics between configurations, and is thus broadly applicable to any sampling technique. Due to the generality of this sampling task we therefore believe that it is similarly suited to benchmark enhanced sampling techniques for various different complex systems.
@@ III.1 Swap Monte Carlo (SMC)
A major problem for benchmarking enhanced sampling techniques is usually the absence of a reference solution and therefore of a clear performance measure. Here, this issue is easily settled by using swap Monte Carlo (SMC) [13, 62, 58]. The algorithm adds non-local Monte Carlo moves on top of conventional molecular dynamics simulations [63]. The Monte Carlo moves are swap moves, in which a pair of particles with different types are randomly selected and their radii are swapped. This swap move is accepted according to a Metropolis scheme. This algorithm is extremely efficient and can be used to equilibrate supercooled liquids at extremely low temperatures, including below the experimental glass transition temperature [58, 60, 56].
@@
The corresponding long-time averages of $\langle E_{\text{pot}}\rangle$ and $c_{V}$ are shown in Fig. 1. We observe that the potential energy decays monotonically for decreasing temperature. Furthermore it can be seen that above $T>T_{\text{SMC}}$ the results for $c_{V}$ are monotonically increasing with decreasing temperature. Once the system falls out of equilibrium near $T_{\rm SMC}\approx 0.12$, there is a strong decrease of $c_{V}$, as found previously in many experiments and simulations in cases where the system falls out of equilibrium [64].
@@ III.2 Molecular dynamics (MD)
Molecular dynamics (MD) simulations consist in solving Newton’s equations of motion with an added thermostat to control the temperature [3]. Consequently, the dynamic relaxation proceeds through realistic dynamics. Our simulations use a Nose-Hoover thermostat with relaxation time $\tau_{\text{NH}}=1.0$ and time step of $\Delta t=0.005.$ Identical to SMC we create high temperature configurations at $T_{\rm init}=0.5$ and then quench the temperature to the desired value $T$ to monitor the relaxation of the potential energy towards equilibrium. After thermalization for $t>10^{7}\,\tau$, we investigate $\langle E_{\text{pot}}\rangle_{t}$ and its fluctuations to measure $c_{V}(t)$. In order to perform a fair comparison, we use the same range of timescales and report the same averaged quantities for MD, SMC and all other sampling methods.
@@
While equilibrium dynamics can easily be measured for MD using for instance time correlation functions, we show instead how the energy decay after a quench from $T_{\rm init}=0.5$ and the time dependence of the specific heat measured during equilibrium sampling in Figs. 2(c,d). Compared to SMC, we show a narrower regime of temperatures down to $T=0.205$. In Figs. 2(c) we observe an increasing time scale to reach the correct value of the energy which becomes impossible for $T<T_{\rm MD}$ over the simulated time window. The lack of sampling becomes more severe when considering the specific heat which can only reach its plateau value for $T=0.32$ but not below. Our data confirm that MD is much less efficient than SMC, as expected. More importantly perhaps, since MD follows the physical dynamics of the system, the data in Fig. 2 in fact serve as a benchmark in order to assess how much gain over the physical dynamics any enhanced algorithm can achieve [17].
@@ III.3 Monte Carlo in temperature space: Parallel tempering (PT)
Parallel tempering [20], also known as replica exchange [21, 19], is a popular enhanced sampling technique applied in a wide range of fields, including spin glasses [21, 19], protein folding [65, 66], polymer melts [67], and solid state physics [68]. In the field of glass-forming liquids, it has been used to create equilibrium structures of deeply supercooled liquids [69, 70], characterize point-to-set length scales [71, 72], and analyze the physics of randomly pinned systems [73].
@@
with energy difference $\Delta U=U(x_{j})-U(x_{j\pm 1})$ between the two configurations and inverse temperatures $\beta_{j}=(k_{B}T_{j})^{-1}$. Using this algorithm configurations evolve both by the physical MD dynamics but also by performing a random walk in temperature space. Low-temperature configurations can therefore follow an “easy” relaxation path by being exchanged with replicas from higher temperatures, then evolving faster at these high temperatures and subsequently being exchanged back to the low temperature. This may avoid configurations being stuck for extremely long times in deep minima at low temperatures. Physically, if the basins relevant at low temperature are also sampled at high temperature, PT can become a very efficient method [27].
The most important factor to optimize PT simulations is the choice of the replica temperatures. On the one hand, large temperature differences will significantly reduce the acceptance rate for exchange events and therefore slow down the exploration of the temperature space. On the other hand, a large number of replicas implies a larger computational effort. After trial and error, we finally use $n=8$ replicas with temperatures $T=$ $0.4$, 0.359, 0.32, 0.287,0.256,0.229, 0.205, $0.183$. This choice is rationalized by a significant overlap between energy distributions at neighboring temperatures and therefore a large acceptance rate $\langle P_{\text{acc}}\rangle>0.25$ for all replicas. We also checked that smaller temperature steps do not improve the results. We optimized the maximal and minimal temperatures in the above range to finally settle on this list of $n=8$ replicas. We also choose $N_{\text{ex}}=5000$ which is a reasonable choice between too frequent or too infrequent temperature swaps, but the results are not very sensitive to this specific choice.

### v2

@@ I Introduction
One of the most important methodological revolution in science in the last century is scientific computing Battimelli et al. (2020). Numerical simulations represent a way, complementary to experiments, to study physical systems, thus providing a unique lens on the microscopic mechanisms underpinning macroscopic physical phenomena Battimelli et al. (2020); Frenkel and Smit (2001); Allen and Tildesley (2017); Newman and Barkema (1999).
@@
A major application of numerical simulations, from the very beginning, has been sampling physical configurations at thermal equilibrium Metropolis et al. (1953). This was initially done by using either Monte Carlo Markov chains Metropolis et al. (1953); Hastings (1970) or molecular dynamics Alder et al. (1957); Rahman and Stillinger (1971). Both methods can be viewed as ways to implement some physical dynamics to ergodically explore the configuration space, just as the physical system does. The basic challenge is to run those dynamics long enough to be able to generate a large set of uncorrelated configurations to perform accurate ensemble averages of physical observables Frenkel and Smit (2001).
@@
When the system is characterized by large relaxation times, for instance near phase transitions or in disordered media, sampling can become so challenging that conventional methods may fail Frenkel and Smit (2001); Newman and Barkema (1999). In such cases, the only solution, so far, consists in devising alternative dynamics that ensure equilibrium sampling while being characterised by substantially smaller decorrelation times. Such strategies are described in many standard textbooks on computer simulations and statistical physics Newman and Barkema (1999); Allen and Tildesley (2017); Krauth (2006); Landau and Binder (2021). In the context of off-lattice molecular simulations, we can mention non-local Gazzillo and Pastore (1989); Kranendonk and Frenkel (1991); Grigera and Parisi (2001), lifting Vucelja (2016), or collective Bernard et al. (2009); Krauth (2021); Ghimenti et al. (2024) Monte Carlo algorithms, parallel tempering Marinari and Parisi (1992); Hukushima and Nemoto (1996); Earl and Deem (2005); Swendsen and Wang (1986), population annealing Hukushima and Iba (2003); Machta (2010); Amey and Machta (2018a), or irreversible Langevin dynamics Ghimenti et al. (2023).
@@
However, there exist physical systems in which equilibrium sampling remains a difficult open problem even with enhanced sampling methods. Among them, glassy systems Berthier and Biroli (2011) stand as one of the most difficult to simulate in condensed and soft-matter physics. Molecular, colloidal and spin glasses are in fact known to display an extremely slow physical dynamics, which creates a major challenge to standard simulation algorithms Berthier and Reichman (2023); Barrat and Berthier (2023). Glassy systems can in fact serve as a severe test of any newly proposed method, and can be seen as a paradigm for complex systems.
@@
The recent discovery of generative models in artificial intelligence (AI) able to generate large structured data such as images, sound, 3D-video, and text has the potential to induce a second revolution in scientific computing Goodfellow et al. (2014); Sohl-Dickstein et al. (2015); Ho et al. (2020); Kingma and Welling (2013); Rezende and Mohamed (2015). These AI models are not only able to accurately produce complex data, but are also very fast. Speed is a central requirement in the realm of scientific computing. Several applications appeared already. In 2019, Noé et al. proposed the usage of normalizing flows (NF) Noé et al. (2019), and independently Wu et al. variational autoregressive models Wu et al. (2019), for Boltzmann sampling in statistical and condensed matter physics. These works have found numerous applications for sampling Invernizzi et al. (2022); Gabrié et al. (2022); Falkner et al. (2023); Coretti et al. (2022); van Leeuwen et al. (2023) and free energy calculations Ding and Zhang (2020); Wirnsberger et al. (2023). However, despite these interesting premises, a clear view on when, where and how these methods work, and in particular their limitations and efficiency against known algorithms, is currently lacking. For standard phase transitions, promising results have been obtained in Marchand et al. (2022); Singha et al. (2023). However, for hard computational problems such as complex and glassy systems, it is unclear whether they can circumvent the problem of large relaxation times. Positive results have been reported in McNaughton et al. (2020); Scriva et al. (2023) for spin-glasses. On the other hand, theoretical and numerical analysis of mean-field models for structural glasses Ciarella et al. (2023); Ghio et al. (2023), related to other hard problems in computer science Mezard and Montanari (2009), has shown that several generative models do not, and sometimes can not, have good performances. Worse, they sometimes perform less efficiently than conventional algorithms, such as local Monte Carlo Ciarella et al. (2023). These results provide a rather pessimistic view of the potentiality of machine learning (ML) techniques in the field of glassy systems. A last difficulty is that the performance of generative models is generically expected to scale very badly with the size of the system, so that applications to study phase transitions and collective effects in many-body systems appear out of sight.
@@
Given the rapid progress made in ML studies Ronhovde et al. (2012); Schoenholz et al. (2016); Bapst et al. (2020); Paret et al. (2020); Jung et al. (2023a), we feel that there is room for hope and progress. At the moment, there is a clear need of further studies to develop and test generative models in hard computational problems, and benchmark their performances against the ones of existing algorithms. The aim of this work is to perform such analysis for atomistic models of glass-forming liquids Berthier and Reichman (2023). Contrary to Ciarella et al. (2023); Ghio et al. (2023), we study an off-lattice, finite-dimensional glassy model which displays an extremely slow dynamics, associated to a super-Arrhenius evolution of relaxation and sampling times. It is challenging to perform numerical simulations in realistic experimental conditions for this model, as the physical relaxation time increases by more than 14 orders of magnitude towards the experimental glass transition temperature Scalliet et al. (2022).
@@
We focus on a specific two-dimensional glass-forming model that shows conventional signatures of glassy dynamics Jung et al. (2023b, 2024), and represents therefore a relevant and challenging test bench for enhanced sampling methods. At low temperatures, molecular dynamics becomes rapidly unable to perform an equilibrium sampling of the configuration space, even for modest system sizes. Given the challenges mentioned above, we intentionally study a relatively small system size to separate the capabilities of NF to tackle complex landscapes from its potentially problematic scaling with system size.
@@
We optimise a ML technique based on normalizing flows Rezende and Mohamed (2015); Noé et al. (2019), which we carefully benchmark against several advanced techniques introduced in the physics literature, such as parallel tempering Hukushima and Nemoto (1996), population annealing Machta (2010), and swap Monte Carlo Ninarello et al. (2017). By studying and comparing their abilities to produce an ensemble of equilibrated low-temperature configurations, we provide the first quantitative analysis of the performance of ML methods to sample realistic supercooled liquids at low temperatures. Surprisingly, our results demonstrate the great potential of such method which turns out to be much more efficient than conventional molecular dynamics and achieves performances comparable to parallel tempering and population annealing. Finally, we assess current limitations of these new methods and provide guidelines for further studies, in particular to improve the parametrization of the normalizing flow and to extend this technique to larger system sizes.
@@ II.1 A two-dimensional ternary Lennard-Jones mixture
We study a two-dimensional ($d=2$) model introduced and developed in Ref. Jung et al. (2023b). This is a variation of the binary Lennard-Jones mixture introduced long ago by Kob and Andersen Kob and Andersen (1995) in which a third component is introduced to both improve the glass-forming ability (i.e. to prevent easy crystallization) and enable a more efficient use of the swap Monte Carlo algorithm, a strategy proposed in Ninarello et al. (2017); Parmar et al. (2020). We refer to Jung et al. (2023b, 2024) for all details regarding the model parameters and simulation details.
Figure 1: Snapshot of a typical amorphous glassy configuration of the two-dimensional model at temperature $T=0.205.$ Colors indicate different types of particles. The goal of this study is to produce a large number of independent configurations drawn from the Boltzmann distribution in Eq. (1).
We investigate systems with $N=43$ particles using periodic boundary conditions with box length $L=6.0$, in reduced units (see Fig. 1 for a snapshot). The unit of length is $\sigma$, which corresponds to the diameter of the large particles. When using molecular dynamics, the unit of time is the Lennard-Jones timescale $\tau=\sqrt{m\sigma^{2}/\epsilon}$ where $m$ is the particle mass, and $\epsilon$ the interaction strength between large particles. For different algorithms, we express times in units of $\tau$, in order to carefully reflect the actual computational cost of each method.
The relatively small system size is actually comparable to previous studies using normalizing flows for sampling in complex systems Noé et al. (2019); Invernizzi et al. (2022); Coretti et al. (2022). The total dimensionality of the problem is $D=Nd=86$. The main goal of this work is to benchmark the efficiency of normalizing flows in sampling such small glassy systems according to the Boltzmann distribution,
$\rho_{*}(x)=Z_{*}^{-1}\exp(-\beta_{*}U(x)).$ (1)
Here, $Z_{*}$ is a proportionality constant, $\beta_{*}$ is the target inverse temperature and $U(x)=E_{\text{pot}}(x)$ is the total (potential) energy. For each configuration, we measure the total potential energy, $E_{\rm pot}=\sum_{i\neq j}V(r_{ij})$, where $V(r)$ denotes the short-range repulsive pair interaction potential, as defined in the Supp. Mat. I of Ref. Jung et al. (2023b), and $r_{ij}$ the relative distance between particles $i$ and $j$.
Results for a larger system size, $N=172$, are presented in Appendix C. Scaling to larger systems introduces additional challenges that have to be considered separately and are left for future work. While $N=43$ seems a small number of particles, we emphasize that this is large enough to produce glassy dynamics and local structure for this dense fluid that are nearly equivalent to those of much larger systems Heuer (2008) (see also Fig. 1). This implies in particular that equilibrium sampling even for this modest system size is already a difficult computational challenge.
@@ II.2 The specific heat as a sampling task
To solve this problem we analyse the statistics of energy fluctuations measured in an ensemble of configurations. We define the average potential energy over this ensemble, $\langle E_{\rm pot}\rangle$, and the variance of its fluctuations, which is directly connected to the specific heat as Allen and Tildesley (2017)
$c_{V}=\frac{C_{V}}{N}=\frac{\langle E_{\rm pot}^{2}\rangle-\langle E_{\rm pot}\rangle^{2}}{Nk_{B}T^{2}}.$ (2)
Here, $T$ denotes the temperature of the system and $k_{B}$ the Boltzmann constant ($k_{B}=1$ in our units). A correct estimate of the specific heat at a given temperature thus requires the production of several independent equilibrium configurations in order to correctly assess the fluctuations around the mean $\langle E_{\rm pot}\rangle$. We conclude that the determination of the specific heat represents a well-defined task that is able to probe the capability of a given algorithm to (i) reach thermal equilibrium, (ii) sample a large number of independent configurations $x$ representative of the Boltzmann distribution. Another advantage is that $c_{V}$ does not require knowledge of a dynamics between configurations, and is thus broadly applicable to any sampling technique. Due to the generality of this sampling task we therefore believe that it is similarly suited to benchmark enhanced sampling techniques for various different complex systems.
@@ III.1 Swap Monte Carlo (SMC)
A major problem for benchmarking enhanced sampling techniques is usually the absence of a reference solution and therefore of a clear performance measure. Here, this issue is easily settled by using swap Monte Carlo (SMC) Grigera and Parisi (2001); Berthier et al. (2016a); Ninarello et al. (2017). The algorithm adds non-local Monte Carlo moves on top of conventional molecular dynamics simulations Berthier et al. (2019). The Monte Carlo moves are swap moves, in which a pair of particles with different types are randomly selected and their radii are swapped. This swap move is accepted according to a Metropolis scheme. This algorithm is extremely efficient and can be used to equilibrate supercooled liquids at extremely low temperatures, including below the experimental glass transition temperature Ninarello et al. (2017); Parmar et al. (2020); Jung et al. (2023b). A detailed introduction and discussion of this algorithm can be found in Ref. Berthier et al. (2019).
@@
The corresponding long-time averages of $\langle E_{\text{pot}}\rangle$ and $c_{V}$ are shown in Fig. 2. We observe that the potential energy decays monotonically for decreasing temperature. Furthermore it can be seen that above $T>T_{\text{SMC}}$ the results for $c_{V}$ are monotonically increasing with decreasing temperature. Once the system falls out of equilibrium near $T_{\rm SMC}\approx 0.12$, there is a strong decrease of $c_{V}$, as found previously in many experiments and simulations in cases where the system falls out of equilibrium Flenner and Szamel (2006).
@@ III.2 Molecular dynamics (MD)
Molecular dynamics (MD) simulations consist in solving Newton’s equations of motion with an added thermostat to control the temperature Allen and Tildesley (2017). Consequently, the dynamic relaxation proceeds through realistic dynamics. Our simulations use a Nose-Hoover thermostat with relaxation time $\tau_{\text{NH}}=1.0$ and time step of $\Delta t=0.005.$ Identical to SMC we create high temperature configurations at $T_{\rm init}=0.5$ and then quench the temperature to the desired value $T$ to monitor the relaxation of the potential energy towards equilibrium. After thermalization for $t>10^{7}\,\tau$, we investigate $\langle E_{\text{pot}}\rangle_{t}$ and its fluctuations to measure $c_{V}(t)$. In order to perform a fair comparison, we use the same range of timescales and report the same averaged quantities for MD, SMC and all other sampling methods.
@@
While equilibrium dynamics can easily be measured for MD using for instance time correlation functions, we show instead how the energy decay after a quench from $T_{\rm init}=0.5$ and the time dependence of the specific heat measured during equilibrium sampling in Figs. 3(c,d). Compared to SMC, we show a narrower regime of temperatures down to $T=0.205$. In Figs. 3(c) we observe an increasing time scale to reach the correct value of the energy which becomes impossible for $T<T_{\rm MD}$ over the simulated time window. The lack of sampling becomes more severe when considering the specific heat which can only reach its plateau value for $T=0.32$ but not below. Our data confirm that MD is much less efficient than SMC, as expected. More importantly perhaps, since MD follows the physical dynamics of the system, the data in Fig. 3 in fact serve as a benchmark in order to assess how much gain over the physical dynamics any enhanced algorithm can achieve Ghimenti et al. (2024).
@@ III.3 Monte Carlo in temperature space: Parallel tempering (PT)
Parallel tempering Earl and Deem (2005), also known as replica exchange Swendsen and Wang (1986); Hukushima and Nemoto (1996), is a popular enhanced sampling technique applied in a wide range of fields, including spin glasses Swendsen and Wang (1986); Hukushima and Nemoto (1996), protein folding Sugita and Okamoto (1999); Bussi et al. (2006), polymer melts Bunker and Dünweg (2000), and solid state physics Falcioni and Deem (1999). In the field of glass-forming liquids, it has been used to create equilibrium structures of deeply supercooled liquids Yamamoto and Kob (2000); De Michele and Sciortino (2002), characterize point-to-set length scales Yaida et al. (2016); Berthier et al. (2016b), and analyze the physics of randomly pinned systems Kob and Berthier (2013).
@@
with energy difference $\Delta U=E_{\text{pot}}(x_{j})-E_{\text{pot}}(x_{j\pm 1})$ between the two configurations and inverse temperatures $\beta_{j}=(k_{B}T_{j})^{-1}$. An extended derivation of this equation and efficient implementation can be found in Chapter 14.1 of Ref. Frenkel and Smit (2001). Using this algorithm configurations evolve both by the physical MD dynamics but also by performing a random walk in temperature space. Low-temperature configurations can therefore follow an “easy” relaxation path by being exchanged with replicas from higher temperatures, then evolving faster at these high temperatures and subsequently being exchanged back to the low temperature. This may avoid configurations being stuck for extremely long times in deep minima at low temperatures. Physically, if the basins relevant at low temperature are also sampled at high temperature, PT can become a very efficient method Berthier and Reichman (2023).
The most important factor to optimize PT simulations is the choice of the replica temperatures. On the one hand, large temperature differences will significantly reduce the acceptance rate for exchange events and therefore slow down the exploration of the temperature space. On the other hand, a large number of replicas implies a larger computational effort. After trial and error, we finally use $n=8$ replicas with temperatures $T=$ $0.4$, 0.359, 0.32, 0.287,0.256,0.229, 0.205, $0.183$. To arrive at this choice, we have started with a small number of replicas and systematically increased their number until we found the optimal result in the given computational time. This choice is rationalized by a significant overlap between energy distributions at neighboring temperatures and therefore a large acceptance rate $\langle P_{\text{acc}}\rangle>0.25$ for all replicas. We also checked that smaller temperature steps do not improve the results. We optimized the maximal and minimal temperatures in the above range to finally settle on this list of $n=8$ replicas. We also choose $N_{\text{ex}}=5000$ which is a reasonable choice between too frequent or too infrequent temperature swaps, but the results are not very sensitive to this specific choice.